OpenStudy (anonymous):

an ellispe has a vertex at (5,0) a co vertex at (0,-3), and center at the origin. which is the equation of the ellispe is standard form?

OpenStudy (anonymous):

is it x^2+y^2/5=1

OpenStudy (amistre64):

the under of x and y should be the square of their "intercepts"

OpenStudy (amistre64):

(5,0) when y = 0, x = 5; so x^2/5^2 = 1 (0,-3) when x = 0, y = -3; so y^2/3^2 = 1